Please help:
For the given functions f and g, find the requested function:
f(x)=2x^2+3; g(x)=3x^2-5x-2
Find (f+g)(3)
Thank you so much!

Answer by stanbon(75887)

For the given functions f and g, find the requested function:
f(x)=2x^2+3; g(x)=3x^2-5x-2
Find (f+g)(3)
------------------------
(f+g)(3) = f(3) + g(3)
---
f(3) = 2*9 + 3 = 21
g(3) = 3*9-15-2 = 10
-----------------
Ans: 31
================================
Cheers,
Stan H.